### v4

@@ 3.1.3 Talking Face Generation
Yao et alyao2021latent ICCV’21 G.+T. By integrating explicit decoupling terms and identity-consistent terms into the loss function, the preservation of facial identity information is improved, resulting in high-quality face editing in videos.
@@
HifaFace gao2021high CVPR’21 G. Proposed a solution based on wavelet transform to address the issue of partial loss of attribute information when generating edited results due to "cyclic consistency" problems.
@@
Preechakul et alpreechakul2022diffusion CVPR’22 D. When encoding images, it is divided into semantically meaningful parts and parts that represent the details of the image.
@@
FENeRF sun2022fenerf CVPR’22 G.+NeRF ➊➍ The introduction of semantic masks into the conditional radiance field enables finer image textures.
@@
GuidedStyle hou2022guidedstyle NN’22 G. Generating faces after face editing is guided based on facial attribute classification. The introduction of a sparse attention mecha -nism enhances the manipulation of individual attribute styles.
@@
FDNeRF zhang2022fdnerf SIG.’22 G.+NeRF The introduction of the Conditional Feature Warping (CFW) module addresses the issue of temporal inconsistency caused by dynamic information in the process of face editing in videos.
@@
AnyFace sun2022anyface CVPR’22 G. ➏➑ Proposed a dual-branch framework for text-driven facial editing, with coordination achieved between the two branches through a Cross-Modal Distillation (CMD) module.
@@
TransEditor xu2022transeditor CVPR’22 G. ➌➊ Emphasizing dual-space GAN interaction’s importance, a transformer architecture is introduced for improved interaction.
@@
Huang et alhuang2023collaborative CVPR’23 D. Proposed the concept of assisted diffusion, integrating individual multimodalities to explore the complementarity between differe -nt modalities.
@@
Ozkan et alozkan2023conceptual ICCV’23 G. The entangled attribute space is decomposed into conceptual and hierarchical latent spaces, and transformer network encoders are employed to modify information in the latent space.
@@
CIPS-3D++ zhou2023cips TPAMI’23 G.+NeRF ➊➒ Replaced the convolutional architecture with an MLP (Multi-Layer Perceptron) architecture to achieve faster rendering speeds.
@@
ClipFace aneja2023clipface SIG.’23 G.+3DMM Learned texture generation from large-scale datasets, enhancing generator performance through generative adversarial training.
@@
TG-3DFace yu2023towards ICCV’23 G. ➌➏ For different scenarios, two sets of text-to-face cross-modal alignment methods were designed with specific focuses.
@@
VecGAN++ dalva2022vecgan TPAMI’23 G. Orthogonal constraint and disentanglement loss are used to decouple attribute vectors in the latent space.
@@
DiffusionRig ding2023diffusionrig CVPR’23 D. 3DMM and diffusion model integration propose a two-stage method for learning personalized facial details.
@@
Kim et alkim2023diffusion CVPR’23 D. Proposed a method for facial editing in videos based on the diffusion model.
@@
SDGAN et alHuang2024SDGAN AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
@@
FaceDNeRF zhang2024facednerf NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.
3.1.4 Facial Attribute Editing
@@
In this section, we review current methods chronologically, following the progression in overcoming technical challenges, primarily focusing on multiple attribute editing methods utilizing GANs. Finally, we summarize methods in Table 5.
@@
$\bullet$ Comprehensive Editing. Facial attribute editing aims to selectively alter specific facial attributes without affecting others. Therefore, disentangling different facial attributes is a primary challenge. Early facial attribute editing models shen2017learning ; zhou2017genegan often achieve editing for a single attribute through data-driven training. For instance, Shen et alshen2017learning propose learning the difference between pre-/post-operation images, represented as residual images, to achieve attribute-specific operations. However, single-attribute editing falls short of meeting expectations, and compression steps in the process often lead to a significant loss of image resolution, a common issue in early methods. The fundamental challenge in comprehensive editing and unrelated attribute modification is achieving complete attribute disentanglement. Many approaches shen2020interpreting ; yao2021latent ; gao2021high ; xu2022transeditor have explored this. E.g., HifaFace gao2021high identifies cycle consistency issues as the cause of facial attribute information loss that proposes a wavelet-based method for high-fidelity face editing, while TransEditor xu2022transeditor introduces a dual-space GAN structure based on the transformer framework that improves image quality and attribute editing flexibility.
@@
$\bullet$ Irrelevant-attribute Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images yao2021latent ; hou2022guidedstyle ; huang2023ia . GuidedStyle hou2022guidedstyle leverages attention mechanisms in StyleGAN karras2019style for the adaptive selection of style modifications for different image layers. IA-FaceS huang2023ia embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches sun2022fenerf ; zhang2022fdnerf ; jiang2022nerffaceediting ; zhou2023cips combine GANs with NeRF mildenhall2020nerf for enhanced spatial awareness capabilities. Specifically, FENeRF sun2022fenerf uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ zhou2023cips enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
@@
$\bullet$ Text Driven facial attribute editing is a crucial application scenario and a recent hot topic in academic research sun2022anyface ; hou2022textface ; aneja2023clipface ; yu2023towards . TextFace hou2022textface introduces text-to-style mapping, directly encoding text descriptions into the latent space of pre-trained StyleGAN. TG-3DFace yu2023towards introduces two text-to-face cross-modal alignment techniques, including global contrastive learning and fine-grained alignment modules, to enhance the high semantic consistency between the generated 3D face and the input text.
$\bullet$ Diffusion-based models have been introduced into facial attribute editing preechakul2022diffusion ; huang2023collaborative ; ding2023diffusionrig ; kim2023diffusion and achieve excellent results. Huang et alhuang2023collaborative propose a collaborative diffusion framework, utilizing multiple pre-trained unimodal diffusion models together for multimodal face generation and editing. DiffusionRig ding2023diffusionrig conditions the initial 3D face model, which helps preserve facial identity information during personalized editing of facial appearance based on the general facial details before the dataset.
@@
Table 6: Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD.
@@
Space Domain Gram-Net liu2020global CVPR’20 ➆➈ ➆➈ The method posits that genuine faces and fake faces exhibit inconsistencies in texture details.
@@
Face X-ray li2020face CVPR’20 ➀➁➂➉ Focusing on boundary artifacts of face fusion for forgery detection.
@@
Zhao et alzhao2021multi CVPR’21 ➀➁➂ A texture enhancement module, an attention generation module, and a bilinear attention pooling mod -ule are proposed to focus on texture details.
@@
Nirkin et alnirkin2021deepfake TPAMI’21 ➀➁➂ Detecting swapped faces by comparing the facial region with its context (non-facial area).
@@
SBIs shiohara2022detecting CVPR’22 ➁➂➇➉ The belief that the more difficult to detect forged faces typically contain more generalized traces of forg -ery can encourage the model to learn a feature representation with greater generalization ability.
RECCE cao2022end CVPR’22 ➀➁➂zi2020wilddeepfake Reconstruction learning on real samples to learn common compressed representations of real images.
@@
LGrad tan2023learning CVPR’23 The gradient is utilized to present generalized artifacts that are fed into the classifier to determine the truth of the image.
@@
NoiseDF wang2023noise AAAI’23 ➀➁➂➃ Extracting noise traces and features from cropped faces and background squares in video frames.
@@
Ba et alba2024exposing AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on information bottleneck theory, is proposed for constraint.
Time Domain Yang et alyang2019exposing ICASSP’19 Focusing on the inconsistency in the head pose in videos by comparing the estimated head pose using all facial landmarks with the one estimated using only the landmarks in the central region.
@@
FTCN zheng2021exploring ICCV’21 ➀➁➂➃li2019faceshifter It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
@@
LipForensics haliassos2021lips CVPR’21 ➀➁➂ Concern about temporal inconsistency of mouth movements in videos.
@@
M2TR wang2022m2tr ICMR’22 ➀➁➂➉ Capturing local inconsistencies at different scales for forgery detection using a multiscale transformer.
@@
Gu et algu2022delving AAAI’22 ➀➁➂zi2020wilddeepfake By densely sampling adjacent frames to pay attention to the inter-frame image inconsistency.
@@
Yang et alyang2023masked TIFS’23 ➀➁➂ ➀➁➂ Treating detection as a graph classification problem and focusing on the relationship between the local image features across different frames.
@@
AVoiD-DF yang2023avoid TIFS’23 ➁➄khalid2021fakeavceleb ➁➄khalid2021fakeavceleb Multimodal forgery detection using audiovisual inconsistency.
@@
Choi et alchoi2024exploiting CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
@@
Xu et alxu2024towards IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
@@
Peng et alpeng2024deepfakes TIFS’24 ➀➂➇ ➀➂➇ Focuse on inter-frame gaze angles, extracting gaze informations and employing spatio-temporal feature aggregation to combine temporal, spatial, and texture features for detection and classification.
Frequency F3-Net qian2020thinking ECCV’20 A two-branch frequency perception framework with a cross-attention module is proposed.
@@
FDFL li2021frequency CVPR’21 Propose an adaptive frequency feature generation module to extract differential features from different frequency bands in a learnable manner.
@@
HFI-Net miao2022hierarchical TIFS’22 ➁➂➃➅korshunov2018deepfakes Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
@@
Guo et alguo2023constructing TIFS’23 ➀➁ ➀➁➂ Designing a backbone network for Deepfake detection with space-frequency interaction convolution.
@@
Tan et altan2024frequency AAAI’24 ➀➄ A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.
@@
Data Driven Dang et aldang2020detection CVPR’20 ➂➅ Utilizing attention mechanisms to handle the feature maps of the detection model.
@@
Zhao et alzhao2021learning ICCV’21 ➀➁➂➃➇➉ Proposes pairwise self-consistent learning for training CNN to extract these source features and detect deep vacation images.
@@
Finfer hu2022finfer AAAI’22 ➀➂➇zi2020wilddeepfake Based on an autoregressive model, using the facial representation of the current frame to predict the facial representation of future frames.
@@
Huang et alhuang2023implicit CVPR’23 ➀➁➂➉li2019faceshifter A new implicit identity-driven face exchange detection framework is proposed.
@@
HiFi-Net guo2023hierarchical CVPR’23 Converting forgery detection and localization into a hierarchical fine-grained classification problem.
@@
Zhai et alzhai2023towards ICCV’23 dong2013casia hsu2006detecting wen2016coverage Weakly supervised image processing detection is proposed such that only binary image level labels (real or tampered) are required for training.
3.2 Forgery Detection

### v5

@@ 3.1.3. Talking Face Generation
Yao et al(yao2021latent, ) ICCV’21 G.+T. By integrating explicit decoupling terms and identity-consistent terms into the loss function, the preservation of facial identity information is improved, resulting in high-quality face editing in videos.
@@
HifaFace (gao2021high, ) CVPR’21 G. Proposed a solution based on wavelet transform to address the issue of partial loss of attribute information when generating edited results due to "cyclic consistency" problems.
@@
Preechakul et al(preechakul2022diffusion, ) CVPR’22 D. When encoding images, it is divided into semantically meaningful parts and parts that represent the details of the image.
@@
FENeRF (sun2022fenerf, ) CVPR’22 G.+NeRF ➊➍ The introduction of semantic masks into the conditional radiance field enables finer image textures.
@@
GuidedStyle (hou2022guidedstyle, ) NN’22 G. Generating faces after face editing is guided based on facial attribute classification. The introduction of a sparse attention mecha -nism enhances the manipulation of individual attribute styles.
@@
FDNeRF (zhang2022fdnerf, ) SIG.’22 G.+NeRF The introduction of the Conditional Feature Warping (CFW) module addresses the issue of temporal inconsistency caused by dynamic information in the process of face editing in videos.
@@
AnyFace (sun2022anyface, ) CVPR’22 G. ➏➑ Proposed a dual-branch framework for text-driven facial editing, with coordination achieved between the two branches through a Cross-Modal Distillation (CMD) module.
@@
TransEditor (xu2022transeditor, ) CVPR’22 G. ➌➊ Emphasizing dual-space GAN interaction’s importance, a transformer architecture is introduced for improved interaction.
@@
Huang et al(huang2023collaborative, ) CVPR’23 D. Proposed the concept of assisted diffusion, integrating individual multimodalities to explore the complementarity between differe -nt modalities.
@@
Ozkan et al(ozkan2023conceptual, ) ICCV’23 G. The entangled attribute space is decomposed into conceptual and hierarchical latent spaces, and transformer network encoders are employed to modify information in the latent space.
@@
CIPS-3D++ (zhou2023cips, ) TPAMI’23 G.+NeRF ➊➒ Replaced the convolutional architecture with an MLP (Multi-Layer Perceptron) architecture to achieve faster rendering speeds.
@@
ClipFace (aneja2023clipface, ) SIG.’23 G.+3DMM Learned texture generation from large-scale datasets, enhancing generator performance through generative adversarial training.
@@
TG-3DFace (yu2023towards, ) ICCV’23 G. ➌➏ For different scenarios, two sets of text-to-face cross-modal alignment methods were designed with specific focuses.
@@
VecGAN++ (dalva2022vecgan, ) TPAMI’23 G. Orthogonal constraint and disentanglement loss are used to decouple attribute vectors in the latent space.
@@
DiffusionRig (ding2023diffusionrig, ) CVPR’23 D. 3DMM and diffusion model integration propose a two-stage method for learning personalized facial details.
@@
Kim et al(kim2023diffusion, ) CVPR’23 D. Proposed a method for facial editing in videos based on the diffusion model.
@@
SDGAN et al(Huang2024SDGAN, ) AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
@@
FaceDNeRF (zhang2024facednerf, ) NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.
NeRFFaceEditing (jiang2025towards, ) TPAMI’25 G.+NeRF Disentangling geometry and appearance within a tri-plane NeRF through statistical tri-plane features and 3D semantic masks.
M-LMPF (zeng2025secure, ) PR’25 G. M-LMPF achieving controllable attribute modification with strong privacy protection and superior editing fidelity.
Some method (chen2024gstalker, ; li2024talkinggaussian, ) introduce 3DGS to achieve more refined facial reconstruction and motion details, aiming to address the issue of insufficient pose and expression control caused by NeRF’s implicit representation.
3.1.4. Facial Attribute Editing
@@
In this section, we review current methods chronologically, following the progression in overcoming technical challenges. Finally, we summarize methods in Tab. 5.
@@
$\bullet$ Comprehensive Editing. Early facial attribute editing models (shen2017learning, ; zhou2017genegan, ) often achieve editing for a single attribute through data-driven training. For instance, Shen et al(shen2017learning, ) propose learning the difference between pre-/post-operation images, represented as residual images, to achieve attribute-specific operations. However, single-attribute editing falls short of meeting expectations, and compression steps in the process often lead to a significant loss of image resolution, a common issue in early methods. The fundamental challenge in comprehensive editing and unrelated attribute modification is achieving complete attribute disentanglement. Many approaches (shen2020interpreting, ; yao2021latent, ; gao2021high, ; xu2022transeditor, ) have explored this. E.g., HifaFace (gao2021high, ) identifies cycle consistency issues as the cause of facial attribute information loss that proposes a wavelet-based method for high-fidelity face editing, while TransEditor (xu2022transeditor, ) introduces a dual-space GAN structure based on the transformer framework that improves image quality and attribute editing flexibility.
@@
$\bullet$ Irrelevant-attribute Retained. Another critical aspect of face editing is retaining as much target image information as possible in the generated images (yao2021latent, ; hou2022guidedstyle, ; huang2023ia, ). GuidedStyle (hou2022guidedstyle, ) leverages attention mechanisms in StyleGAN (karras2019style, ) for the adaptive selection of style modifications for different image layers. IA-FaceS (huang2023ia, ) embeds the face image to be edited into two branches of the model, where one branch calculates high-dimensional component-invariant content embedding to capture facial details, and the other branch provides low-dimensional component-specific embedding for component operations. Additionally, some approaches (sun2022fenerf, ; zhang2022fdnerf, ; jiang2022nerffaceediting, ; zhou2023cips, ) combine GANs with NeRF (mildenhall2020nerf, ) for enhanced spatial awareness capabilities. Specifically, FENeRF (sun2022fenerf, ) uses two decoupled latent codes to generate corresponding facial semantics and textures in a 3D volume with spatial alignment sharing the same geometry. CIPS-3D++ (zhou2023cips, ) enhances the model’s training efficiency with a NeRF-based shallow 3D shape encoder and an MLP-based deep 2D image decoder.
@@
$\bullet$ Text Driven. Text driven facial attribute editing is a crucial application scenario and a recent hot topic in academic research (sun2022anyface, ; hou2022textface, ; aneja2023clipface, ; yu2023towards, ). TextFace (hou2022textface, ) introduces text-to-style mapping, directly encoding text descriptions into the latent space of pre-trained StyleGAN.
@@
Table 6. Overview of representative forgery detection methods. Notations: ➀ FF++, ➁ DFDC, ➂ Celeb-DF, ➃ Deeperforensics, ➄ Self-build, ➅ UADFV, ➆ Celeb-HQ, ➇ DFDCp, ➈ FFHQ, ➉ DFD.
@@
Space Domain Gram-Net (liu2020global, ) CVPR’20 ➆➈ ➆➈ The method posits that genuine faces and fake faces exhibit inconsistencies in texture details.
@@
Face X-ray (li2020face, ) CVPR’20 ➀➁➂➉ Focusing on boundary artifacts of face fusion for forgery detection.
@@
Zhao et al(zhao2021multi, ) CVPR’21 ➀➁➂ A texture enhancement module, an attention generation module, and a bilinear attention pooling mod -ule are proposed to focus on texture details.
@@
Nirkin et al(nirkin2021deepfake, ) TPAMI’21 ➀➁➂ Detecting swapped faces by comparing the facial region with its context (non-facial area).
@@
SBIs (shiohara2022detecting, ) CVPR’22 ➁➂➇➉ The belief that the more difficult to detect forged faces typically contain more generalized traces of forg -ery can encourage the model to learn a feature representation with greater generalization ability.
@@
LGrad (tan2023learning, ) CVPR’23 The gradient is utilized to present generalized artifacts that are fed into the classifier to determine the truth of the image.
@@
NoiseDF (wang2023noise, ) AAAI’23 ➀➁➂➃ Extracting noise traces and features from cropped faces and background squares in video frames.
@@
Ba et al(ba2024exposing, ) AAAI’24 ➀➁➂ ➀➁➂ Multiple non-overlapping local representations are extracted from the image for forgery detection. A local information loss function, based on information bottleneck theory, is proposed for constraint.
UNITE (kundu2025towards, ) CVPR’25 ➀➂➃➅ ➀➂➃➅ Leveraging domain-agnostic features, attention-diversity regularization, and mixed-domain training, achieving great performance on both facial and fully synthetic manipulations.
@@
Time Domain FTCN (zheng2021exploring, ) ICCV’21 ➀➁➂➃(li2019faceshifter, ) It is believed that most face video forgeries are generated frame by frame. As each altered face is inde -pendently generated, this inevitably leads to noticeable flickering and discontinuity.
@@
LipForensics (haliassos2021lips, ) CVPR’21 ➀➁➂ Concern about temporal inconsistency of mouth movements in videos.
@@
M2TR (wang2022m2tr, ) ICMR’22 ➀➁➂➉ Capturing local inconsistencies at different scales for forgery detection using a multiscale transformer.
@@
Gu et al(gu2022delving, ) AAAI’22 ➀➁➂(zi2020wilddeepfake, ) By densely sampling adjacent frames to pay attention to the inter-frame image inconsistency.
@@
Yang et al(yang2023masked, ) TIFS’23 ➀➁➂ ➀➁➂ Treating detection as a graph classification problem and focusing on the relationship between the local image features across different frames.
@@
AVoiD-DF (yang2023avoid, ) TIFS’23 ➁➄(khalid2021fakeavceleb, ) ➁➄(khalid2021fakeavceleb, ) Multimodal forgery detection using audiovisual inconsistency.
@@
Choi et al(choi2024exploiting, ) CVPR’24 ➀➂➃ ➀➂ Focus on the inconsistency of the style latent vectors between frames.
@@
Xu et al(xu2024towards, ) IJCV’24 ➀➁➂➃ ➀➁➂➃ Forgery detection is conducted by converting video clips into thumbnails containing both spatial and temporal information.
@@
Peng et al(peng2024deepfakes, ) TIFS’24 ➀➂➇ ➀➂➇ Focuse on inter-frame gaze angles, extracting gaze informations and employing spatio-temporal feature aggregation to combine temporal, spatial, and texture features for detection and classification.
@@
Frequency FDFL (li2021frequency, ) CVPR’21 Propose an adaptive frequency feature generation module to extract differential features from different frequency bands in a learnable manner.
@@
HFI-Net (miao2022hierarchical, ) TIFS’22 ➁➂➃➅(korshunov2018deepfakes, ) Notice that the forgery flaws used to distinguish between real and fake faces are concentrated in the mid- and high-frequency spectrum.
@@
Guo et al(guo2023constructing, ) TIFS’23 ➀➁ ➀➁➂ Designing a backbone network for Deepfake detection with space-frequency interaction convolution.
@@
Tan et al(tan2024frequency, ) AAAI’24 ➀➄ A lightweight frequency-domain learning network is proposed to constrain classifier operation within the frequency domain.
WMamba (peng2025wmamba, ) MM’25 ➁➂➇ Introduces a Mamba-based wavelet feature extractor that leverages dynamic contour convolution and efficient long-range modeling to capture fine-grained, globally distributed forgery cues.
WaveDIF (dutta2025wavedif, ) CVPR’25 ➀➂ ➀➂ A lightweight frequency-domain detector that leverages DFT-based denoising and wavelet subband en -ergy analysis to achieve competitive accuracy on both intra- and cross-dataset evaluations.
@@
Data Driven Dang et al(dang2020detection, ) CVPR’20 ➂➅ Utilizing attention mechanisms to handle the feature maps of the detection model.
@@
Zhao et al(zhao2021learning, ) ICCV’21 ➀➁➂➃➇➉ Proposes pairwise self-consistent learning for training CNN to extract these source features and detect deep vacation images.
@@
Finfer (hu2022finfer, ) AAAI’22 ➀➂➇(zi2020wilddeepfake, ) Based on an autoregressive model, using the facial representation of the current frame to predict the facial representation of future frames.
@@
Huang et al(huang2023implicit, ) CVPR’23 ➀➁➂➉(li2019faceshifter, ) A new implicit identity-driven face exchange detection framework is proposed.
@@
HiFi-Net (guo2023hierarchical, ) CVPR’23 Converting forgery detection and localization into a hierarchical fine-grained classification problem.
@@
Zhai et al(zhai2023towards, ) ICCV’23 (dong2013casia, ) (hsu2006detecting, )(wen2016coverage, ) Weakly supervised image processing detection is proposed such that only binary image level labels (real or tampered) are required for training.
TG-3DFace (yu2023towards, ) introduces two text-to-face cross-modal alignment techniques, including global contrastive learning and fine-grained alignment modules, to enhance the high semantic consistency.
$\bullet$ Diffusion-based. Diffusion-based models have been introduced into facial attribute editing (preechakul2022diffusion, ; huang2023collaborative, ; ding2023diffusionrig, ; kim2023diffusion, ) and achieve excellent results. Huang et al(huang2023collaborative, ) propose a collaborative diffusion framework, utilizing multiple pre-trained unimodal diffusion models together for multimodal face generation and editing. DiffusionRig (ding2023diffusionrig, ) conditions the initial 3D face model, which helps preserve facial identity information during personalized editing of facial appearance.
3.2. Forgery Detection